Project

General

Profile

2018-01-09-MEMO/MIPO

Joint MICE Experiment Management Office and MICE International Project Office

9th January 2017; 15:30 GMT
RAL, R1 2nd Floor East, MICE Meeting room

Phone details

Notes:

Present: AK, KL, CW, PG, JB, PS, ABr, CR, MP

  1. Introduction: KL
    • Welcome to first combined MEMO/MIPO of 2018. Experiment is entering a new phase with pressure to publish com=bined with decommissioning and making safe. KL and CW propose now to run combined MEMO/MIPO.
  2. Notes of previous meeting and status of actions: All
    • Notes posted on the wiki were accepted. The status of the outstanding actions is:
      • KL: Send note to A.Kurup to alert him to warnings that need to be addressed before on 24Nov17.
        • Done.
      • CR: Define settings for next block data to MIPO/MEMO for Monday 27Nov17.
        • Done.
      • CW: prepare plan for post-run (Cycle 2017/03) survey.
        • Agenda item.
  3. Memo on measurements required before experiment is decommissioned: CR
    • CR guided us through the note posted on the wiki (2017-12-06_decommissioning-measurements_v3.pdf). Points noted:
      • The QD/QP systems had been powered off over the Christmas break. To ensure that magnets can be run, the QD/QP systems must be reinstated as soon as possible. It was agreed that CW would do this as a matter of urgency.
      • It was agreed that CW would consut and prepare a plan to carry out a limitted magnetic-field survey.
      • CR will maintain the `requirements' memo as a living document. We noted the pressure on local resources required to carry out the mechanical and magnetic survey.
  4. Magnet status; stability in quiescent state: JB
    • Focus coil is running at a little higher pressure because of the slight increase in the ambient temperature. This can be recovered.
    • Compressors are leaking helium. May need to presurise the lines once more.
    • Overall the systems are presently running OK.
  5. Hall status and stability of plant and other equipment in quiescent state: CW
    • Systems are more-or-less OK. No big issues to be dealt with though several minor issues need to be addressed (e.g. beam stop, ODH alarms, crane certification etc.).
  6. Planning for decommissioning and making safe to be complete by 31Mar18: CW
    • It was noted that a formal decommissioning plan needs to be presented for endorsement at the MICE/ISIS Safety meeting which will take place at the end of January. The plan must be completed by the end of March 2018. This omplies starting earlier; CW's present estimate is that this will require starting at the beginning of March 2018.
    • US personnel plan to come to RAL in the first week of March 2018 to begin the decommissioning of the components of the experiment that are their responsibility.
  7. Software and computing status: DR
    • No report as DR is travelling back to RAL.
  8. Data validation ahead of start of Cycle 2017/04: CR
    • CR has started to review a representative sample of all data sets to see whether there are issues popping up. Detector experts will be looking at the data too and will present their status at the VC. We noted the time pressure on making a decision to run and the need to make a robust case to generate the necessary permissions to run.
  9. Paper status and plans: CR
    • Emittance paper: Paper has been updated and a wise-person meeting will take place next Monday, 15Jan18;
    • LiH scattering paper: progress has been made on implementing an improved fiducial cut. New version for analysis w/s 18Jan18. Noted that the diffuser cut is presently missing in the analysis. This will be implemented.
    • Emittance reduction paper: data/MC agreement has been improved. Work continues.
    • Routines and procedures have been prepared to standardisation of cuts, calculation of systematic uncertainties and presentation of resutls; all to streamline the analysis and preparation of data for publication. CR/KL will continue to drive this through.
    • Analysis w/s 18Jan18 at RAL
    • CM50 will be preceded by a writing w/s for the papers in progress.
    • Analysis w/s 12/13 April Ggo.
    • Need now to have h/w papers in this list too (CW). KL/CW will meet to make a step in this process.
  10. AOB
    • We agreed to meet fortnightly.

Summary of actions:

  • CW: Reinstate the QD/QP systems.
  • CW: Prepare plan to carry out a magnetic-field survey for discussion.
  • CR: Maintain the 'requirements' document as necessary.
  • KL/CW: Make a step in the preparation of the h/w papers.

Back to MEMO:

Back

Back to MIPO:

Back

Documents

2017-12-06_decommissioning-measurements_v3.pdf (2.59 MB) Rogers, Chris, 09 January 2018 15:32